# Integrable deformations of dimensionally reduced gravity

### v1

@@ 2 Dimensionally reduced gravity
Given a symmetric space $G/K$ the global isometry $G$ has a Lie algebra $\mathfrak{g}$ that decomposes as
@@
where tildes denote the Matzner-Misner scalar fields, to remark their difference with the Ehlers ones. The Lagrangians (2.19), (2.20) must be supplemented with their associated Virasoro constraints, formally obtained by varying the actions with respect to the unimodular degrees of freedom before gauge-fixing,
$V_{(\text{MM})\;\mu\nu}=\partial_{\mu}\rho\partial_{\nu}\tilde{\sigma}-\frac{1}{2}\partial_{\nu}\partial_{\mu}\rho-\frac{1}{2}\rho\text{Tr}(\tilde{P}_{\mu}\tilde{P}_{\nu})=0\,,$ (2.21)
$V_{(\text{E})\;\mu\nu}=\partial_{\mu}\rho\partial_{\nu}\sigma-\frac{1}{2}\partial_{\nu}\partial_{\mu}\rho-\frac{1}{2}\rho\text{Tr}(P_{\mu}P_{\nu})=0\,.$ (2.22)
@@ 2.1.2 A centrally extended Lax connection
However, we here want to stress that, although corresponding to the field equations through its flatness, not only is the Lax connection (2.37) oblivious to the Virasoro constraints (2.22), but, due to the presence of the central charge, it is not even invariant under $\tau^{\infty}$ anymore. Therefore, the extent of its usefulness in the construction of the monodromy matrix, which is usually spacetime independent precisely because of such invariance, remains unclear.
@@ 2.2 Hamiltonian integrability
$\{\pi^{(\rho)}(x),\rho(y)\})$ =$=\{\pi^{(\sigma)}(x),\sigma(y)\}=-\delta(x-y)$
$\{\pi_{\mathbf{1}}^{(P)}(x),\pi_{\mathbf{2}}^{(P)}(y)\}$ $=[\Omega_{\mathfrak{p}},\pi^{(Q)}_{\mathbf{2}}]\delta(x-y)\approx 0$
$\{\pi_{\mathbf{1}}^{(P)}(x),\pi_{\mathbf{2}}^{(Q)}(y)\}$ $=[\Omega_{\mathfrak{k}},\pi^{(P)}_{\mathbf{2}}]\delta(x-y)$
$\{\pi_{\mathbf{1}}^{(Q)}(x),\pi_{\mathbf{2}}^{(Q)}(y)\}$ $=[\Omega_{\mathfrak{k}},\pi^{(Q)}_{\mathbf{2}}]\delta(x-y)\approx 0$
$\{\pi_{\mathbf{1}}^{(P)}(x),P_{x,\mathbf{2}}(y)\}$ $=[\Omega_{\mathfrak{p}},Q_{x,\mathbf{2}}]\delta(x-y)+\Omega_{\mathfrak{p}}\partial_{x}\delta(x-y)$
$\{\pi_{\mathbf{1}}^{(Q)}(x),Q_{x,\mathbf{2}}(y)\}$ $=[\Omega_{\mathfrak{p}},Q_{x,\mathbf{2}}]\delta(x-y)+\Omega_{\mathfrak{k}}\partial_{x}\delta(x-y)$
We introduced $\pi^{(Q)}$ as the canonical variable $\pi^{(Q)}\approx 0$ is the constraint that generates the $K$-gauge symmetry. We used the standard notation $V_{\mathbf{1}}=T\otimes 1$, $V_{\mathbf{2}}=1\otimes T$, where $\Omega=\Omega_{\mathfrak{k}}+\Omega_{\mathfrak{p}}=T_{A}\otimes T^{A}$ is the tensorial Casimir on $\mathfrak{g}$ decomposed into the $\mathbb{Z}_{2}$-grading $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$. In these canonical variables the Lax matrix $\mathfrak{L}_{ext,x}$ is
@@
With the classical $r$-matrix444we remark that $L_{0}=-z\frac{\partial}{\partial z}$
@@
${}\quad-\frac{1}{2}\left(\frac{1}{\rho(x)}+\frac{1}{\rho(y)}\right)\left(r_{\mathbf{12}}(z,z^{\prime})+r_{\mathbf{21}}(z^{\prime},z)\right)\partial_{x}\delta(x{-}x^{\prime})\,,$
@@ 3.1 The Lagrangian
which will turn out to be crucial in subsequent computations. For calculations it is often useful to use light-cone coordinates. Then the tensor $K$ is specified as follows in terms of the deformation defining function $E(\nu)$:
@@
$D\left(\rho\star\frac{1+K}{1-K}P\right)$ $=0$ (3.13)
$\mathrm{d}\left(\star\frac{1+K}{1-K}\mathrm{d}\rho\right)$ $=0$ (3.14)
@@ 3.3.2 Virasoro constraints
We show in appendix C.1 that the vanishing of the last entry in (3.24) furnishes the Virasoro constraints (in lightcone coordinates)
@@ 3.3.3 The centrally extended Lax connection
Flatness of (3.29) will be equivalent to the equations of motion (3.5), (3.4) and (3.6), together with the Bianchi identities (2.5), (we adopt form notation in order to avoid cluttering once again)
@@
where we have used (3.30) and the fact that there exists no invariant bilinear pairing between $\mathsf{L}_{0}$ and $T_{A}$.
@@ 3.4 Hamiltonian analysis
In canonical form the Lax matrix takes the form:
$\mathfrak{L}_{x}=Q_{x}+\frac{1+z^{2}}{1-z^{2}}P_{x}+\frac{2z}{1-z^{2}}\frac{\pi^{(P)}}{\rho}+\left(\frac{1+z^{2}}{1-z^{2}}\frac{\partial_{x}\rho}{\rho}+\frac{2z}{1-z^{2}}\frac{\pi^{(\sigma)}}{\rho}\right)L_{0}+\pi^{(\rho)}\mathsf{K}$ (3.35)
@@
As in the other cases of such Auxiliary Field Deformations the canonical analysis of integrability, i.e. the canonical form of the Lax matrix and the involution of the charges, is not affected by the deformation and agrees with the expression in the undeformed case (2.46). This the above argument means that the auxiliary field model has the same classical $r$-matrix in comparison to the undeformed model (2.48).
@@ 4.1 Yang-Baxter deformation of $D=2$ dimensionally reduced gravity
Integrable $\sigma$-models are quite sparse. For a long time, only $\sigma$-models with very particular and symmetric target spaces – flat space, Lie (super)groups, symmetric spaces – have been found to be integrable. A typical route to obtain new integrable models is to deform such known models at the level of the action. One example of non-trivial deformations has been introduced and investigated in the last decade via the so-called Yang-Baxter deformations [23, 24, 25, 64, 65, 66, 67]. These deformations are generated by so-called $R$-operators – algebraic objects based on an underlying Lie algebra structure: solutions to the classical or the modified classical Yang-Baxter equation. In the former case these are typically called homogeneous Yang-Baxter deformation, in the latter $\eta$-deformation. Homogeneous Yang-Baxter deformations have been shown to be closely connected to abelian and non-abelian $T$-duality transformations. [68, 69, 70, 71, 72, 73]. $\eta$-deformations, on the other hand, are non-trivial deformations of the classical dynamics and have novel target space interpretations [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 31] in the case of $\sigma$-models. By now, Yang-Baxter deformations have also been shown to fit into the landscape of classical Gaudin models and 4d Chern-Simons theory [84, 85, 86].
So far, Yang-Baxter deformations have been considered mainly applied to string $\sigma$-models. Here, we introduce the following Lagrangian as the Yang-Baxter deformation of $D=2$ dilaton-gravity coupled to the coset model, (2.20):
@@
for $m,n\in\mathfrak{g}$. The choices $c^{2}=0,1,-1$ correspond to homogeneous, split and non-split case, see [87] for a review. The constant $C$ in (4.2) comes from the identification $\eta=C\rho$, to be justified below, in comparison to the standard form of the Yang-Baxter deformation (for the $\sigma$-model part):
$L=-\frac{\rho}{2}\text{tr}\left(P_{+}\frac{1}{1-\eta\tilde{R}_{g}\circ\mathbb{P}_{\mathfrak{p}}}P_{-}\right)\,.$ (4.3)
@@
    $\eta=0$ this corresponds to the standard undeformed dimensionally reduced gravity.
@@
    $\mathrm{d}\eta=\mathrm{d}\rho=0$, this would correspond to the standard Yang-Baxter deformation of the coset $\sigma$-model.
@@ 4.2 Equations of motion
Remarkably, the second equation (4.6) arises like that from the variation with respect to $\rho$, also in its new role as a deformation parameter in (4.1). This equation is a novel feature in comparison to the cases of the other integrable models in [33], where the ‘running coupling’ $\rho$ was not considered to be a dynamical field.
@@
00 0 $=\mathrm{d}P+\frac{1}{2}[Q,P]=\mathrm{d}K^{(1)}+[K^{(0)},K^{(1)}]$ (4.11)
@@ 4.3 Lax connection
As in [33], we make the same ansatz for the Lax pair as in the case with constant couplings555deformation parameter and constant in front of $\sigma$-model action, namely
$\mathfrak{L}(\lambda)_{Y.B.}=K^{(0)}+\sqrt{\beta(\rho)}\left(\frac{1+\gamma^{2}}{1-\gamma^{2}}K^{(1)}+\frac{2\gamma}{1-\gamma^{2}}\star K^{(1)}\right)$ (4.12)
@@
$\mathrm{d}\star\mathrm{d}\rho-\frac{\beta^{\prime}(\rho)}{\beta(\rho)}\mathrm{d}\rho\wedge\star\mathrm{d}\rho=0.$ (4.17)
@@ 5 Outlook
A key question also originally motivating in part this work was whether it is possible to extend the techniques of integrable $D=2$ field theories to cases with dilaton potentials, i.e.
$\Box\rho=-V(\rho).$ (5.1)
with a scalar potential $V$. Neither of the deformations considered in this article did lead to such a scalar potential: the Auxiliary Field Deformation sources higher derivative modifications (3.20) and the Yang-Baxter deformations lead to a modified equation for the dilaton, which is second order in derivatives (4.17). It would therefore be compelling to understand if other kinds of integrable deformations could lead to such scalar potentials instead. This would provide new opportunities to study cosmologies in $D=2$ with the methods of integrability.
@@ C.1 Virasoro constraint from $\tau^{\infty}$ invariance of the linear system
We here follow the reasoning from [53] in order to derive the Virasoro constraints from the vanishing of the $\mathsf{K}$-coefficient in (3.24). We reformulate such condition in terms of lightcone coordinates
@@
leads to the additional property
@@ D.1 Review: Yang-Baxter deformation of Principal Chiral Model with spacetime dependent couplings
$\mathrm{d}\eta-\eta\frac{\mathrm{d}\rho}{\rho}=0$ (D.6)
@@
$\mathfrak{L}(\lambda)=(1-c^{2}\eta^{2})\left(\frac{1}{1-\lambda^{2}}k+\frac{\lambda}{1-\lambda^{2}}\star k\right)$ (D.8)
@@
$\frac{\mathrm{d}\lambda}{\lambda}=\frac{1}{(1-c^{2}\eta^{2})(1-\lambda^{2})}\left(\left((1+\lambda^{2})(1+c^{2}\eta^{2})-4c^{2}\eta^{2}\right)\frac{\mathrm{d}\eta}{\eta}+\frac{2(\lambda^{2}-c^{2}\eta^{2})}{\lambda}\frac{\star\mathrm{d}\eta}{\eta}\right)$ (D.10)
@@
$\mathrm{d}\star\mathrm{d}\eta-\frac{\beta^{\prime}(\eta)}{\beta(\eta)}\mathrm{d}\eta\wedge\star\mathrm{d}\eta=0.$ (D.12)

### v2

@@ 2 Dimensionally reduced gravity
Given a symmetric space $G/K$, the global isometry $G$ has a Lie algebra $\mathfrak{g}$ that decomposes as
@@
where tildes denote the Matzner-Misner scalar fields, to remark their difference with the Ehlers ones. The Lagrangians (2.19), (2.20) must be supplemented with their associated Virasoro constraints: the latter stem from a combination of the diffeomorphism and the Hamiltonian constraints, formally obtained by varying the actions with respect to the unimodular degrees of freedom before gauge-fixing. In lightcone coordinates (2.7), the Virasoro constraints read
$V_{(\text{MM})\;\pm\pm}=\partial_{\pm}\rho\partial_{\pm}\tilde{\sigma}-\frac{1}{2}\partial_{\pm}\partial_{\pm}\rho-\frac{1}{2}\rho\text{Tr}(\tilde{P}_{\pm}\tilde{P}_{\pm})=0\,,$ (2.21)
$V_{(\text{E})\;\pm\pm}=\partial_{\pm}\rho\partial_{\pm}\sigma-\frac{1}{2}\partial_{\pm}\partial_{\pm}\rho-\frac{1}{2}\rho\text{Tr}(P_{\pm}P_{\pm})=0\,.$ (2.22)
@@ 2.1.2 A centrally extended Lax connection
However, we here want to stress that, although corresponding to the field equations through its flatness, not only is the Lax connection (2.37) oblivious to the Virasoro constraints (2.22), but, due to the presence of the central charge, it is not even invariant under $\tau^{\infty}$ anymore. Therefore, the extent of its usefulness in the construction of the monodromy matrix, which is usually spacetime independent precisely because of such invariance, remains unclear. On the other hand, the present underlying structure strongly resembles the one appearing in Affine Gaudin models [59, 60] – due to appearance of Kac-Moody algebras and central extensions, but not one related to the spectral parameter. [^4] We thank Sylvain Lacroix for pointing this out to us.
@@ 2.2 Hamiltonian integrability
$\{\pi^{(\rho)}(x),\rho(y)\})$ =$=\{\pi^{(\sigma)}(x),\sigma(y)\}=-\delta(x-y),$
$\{\pi_{\mathbf{1}}^{(P)}(x),\pi_{\mathbf{2}}^{(P)}(y)\}$ $=[\Omega_{\mathfrak{p}},\pi^{(Q)}_{\mathbf{2}}]\delta(x-y)\approx 0,$
$\{\pi_{\mathbf{1}}^{(P)}(x),\pi_{\mathbf{2}}^{(Q)}(y)\}$ $=[\Omega_{\mathfrak{k}},\pi^{(P)}_{\mathbf{2}}]\delta(x-y),$
$\{\pi_{\mathbf{1}}^{(Q)}(x),\pi_{\mathbf{2}}^{(Q)}(y)\}$ $=[\Omega_{\mathfrak{k}},\pi^{(Q)}_{\mathbf{2}}]\delta(x-y)\approx 0,$
$\{\pi_{\mathbf{1}}^{(P)}(x),P_{x,\mathbf{2}}(y)\}$ $=[\Omega_{\mathfrak{p}},Q_{x,\mathbf{2}}]\delta(x-y)+\Omega_{\mathfrak{p}}\partial_{x}\delta(x-y),$
$\{\pi_{\mathbf{1}}^{(Q)}(x),Q_{x,\mathbf{2}}(y)\}$ $=[\Omega_{\mathfrak{p}},Q_{x,\mathbf{2}}]\delta(x-y)+\Omega_{\mathfrak{k}}\partial_{x}\delta(x-y).$
We introduced $\pi^{(Q)}$ as the canonical variable $\pi^{(Q)}\approx 0$ is the constraint that generates the $K$-gauge symmetry. We used the standard notation $V_{\mathbf{1}}=V\otimes 1$, $V_{\mathbf{2}}=1\otimes V$, where $\Omega=\Omega_{\mathfrak{k}}+\Omega_{\mathfrak{p}}=T_{A}\otimes T^{A}$ is the tensorial Casimir on $\mathfrak{g}$ decomposed into the $\mathbb{Z}_{2}$-grading $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$. In these canonical variables the Lax matrix $\mathfrak{L}_{ext,x}$ is
@@
With the classical $r$-matrix [^5] We remark that $L_{0}=-z\frac{\partial}{\partial z}$.
@@
${}\quad-\frac{1}{2}\left(\frac{1}{\rho(x)}+\frac{1}{\rho(x^{\prime})}\right)\left(r_{\mathbf{12}}(z,z^{\prime})+r_{\mathbf{21}}(z^{\prime},z)\right)\partial_{x}\delta(x{-}x^{\prime})\,,$
@@ 3.1 The Lagrangian
which will turn out to be crucial in subsequent computations. For calculations it is often useful to use light-cone coordinates (2.7). Then the tensor $K$ is specified as follows in terms of the deformation defining function $E(\nu)$:
@@
$D\left(\rho\star\frac{1+K}{1-K}P\right)$ $=0,$ (3.13)
$\mathrm{d}\left(\star\frac{1+K}{1-K}\mathrm{d}\rho\right)$ $=0,$ (3.14)
@@ 3.3.2 Virasoro constraints
We show in appendix C.1 that the vanishing of the last entry in (3.24) furnishes the Virasoro constraints (in lightcone coordinates (2.7))
@@ 3.3.3 The centrally extended Lax connection
Flatness of (3.29) will be equivalent to the equations of motion (3.5), (3.4) and (3.6), together with the Bianchi identities (2.5) (we adopt form notation in order to avoid cluttering once again),
@@
where we have used (3.30) and the fact that there exists no invariant bilinear pairing between $L_{0}$ and $T_{A}$.
@@ 3.4 Hamiltonian analysis
In Hamiltonian formalism the Lax matrix becomes:
$\mathfrak{L}_{x}=Q_{x}+\frac{1+z^{2}}{1-z^{2}}P_{x}+\frac{2z}{1-z^{2}}\frac{\pi^{(P)}}{\rho}+\left(\frac{1+z^{2}}{1-z^{2}}\frac{\partial_{x}\rho}{\rho}+\frac{2z}{1-z^{2}}\frac{\pi^{(\sigma)}}{\rho}\right)L_{0}+\pi^{(\rho)}\mathsf{K}.$ (3.35)
@@
As in the other cases of such Auxiliary Field Deformations the canonical analysis of integrability, i.e. the canonical form of the Lax matrix and the involution of the charges, is not affected by the deformation and agrees with the expression in the undeformed case (2.46). The above argument means that the auxiliary field model has the same classical $r$-matrix in comparison to the undeformed model (2.48).
@@ 4.1 Yang-Baxter deformation of $D=2$ dimensionally reduced gravity
Integrable $\sigma$-models are quite sparse. For a long time, only $\sigma$-models with very particular and symmetric target spaces – flat space, Lie (super)groups, symmetric spaces – have been found to be integrable. A typical route to obtain new integrable models is to deform such known models at the level of the action. One example of non-trivial deformations has been introduced and investigated in the last decade via the so-called Yang-Baxter deformations [23, 24, 25, 66, 67, 68, 69, 70]. These deformations are generated by so-called $R$-operators – algebraic objects based on an underlying Lie algebra structure: solutions to the (modified) classical Yang-Baxter equation. In the former case these are typically called homogeneous Yang-Baxter deformation, in the latter $\eta$-deformation. Homogeneous Yang-Baxter deformations have been shown to be closely connected to abelian and non-abelian $T$-duality transformations [71, 72, 73, 74, 75, 76]. $\eta$-deformations, on the other hand, are non-trivial deformations of the classical dynamics and have novel target space interpretations [77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 31] in the case of $\sigma$-models. By now, Yang-Baxter deformations have also been shown to fit into the landscape of classical Gaudin models and 4d Chern-Simons theory [87, 60, 88].
So far, Yang-Baxter deformations have been considered mainly applied to string $\sigma$-models. Here, we introduce the following Lagrangian as the Yang-Baxter deformation of $D=2$ dilaton-gravity coupled to the coset $\sigma$-model [^6] Again, for the sake of integrability, we can choose a generic symmetric space $G/K$ as target space. Concrete examples of dimensionally reduced gravity correspond to specific choices of $G/K$, as mentioned in section 2., (2.20):
@@
for $m,n\in\mathfrak{g}$. The choices $c^{2}=0,1,-1$ correspond to homogeneous, split and non-split case, see [89] for a review. In particular, the space of solutions of this equation and, hence, also of possible Yang-Baxter deformations is unchanged in comparison to pure $\sigma$-model case.
The constant $C$ in (4.2) comes from the identification $\eta=C\rho$, to be justified below, in comparison to the standard form of the Yang-Baxter deformation (for the $\sigma$-model part):
$L=-\frac{\rho}{2}\text{tr}\left(P_{+}\frac{1}{1-\eta\tilde{R}_{g}\circ\mathbb{P}_{\mathfrak{p}}}P_{-}\right)\,,$ (4.3)
@@
    $\eta=0$. This corresponds to the standard undeformed dimensionally reduced gravity.
@@
    $\mathrm{d}\eta=\mathrm{d}\rho=0$. This would correspond to the standard Yang-Baxter deformation of the coset $\sigma$-model.
@@ 4.2 Equations of motion
Remarkably, the second equation (4.6) takes this form from the variation with respect to $\rho$, also in its new role as a deformation parameter in (4.1). This equation is a novel feature in comparison to the cases of the other integrable models in [33], where the ‘running coupling’ $\rho$ was not considered to be a dynamical field.
@@
00 0 $=\mathrm{d}P+\frac{1}{2}[Q,P]=\mathrm{d}K^{(1)}+[K^{(0)},K^{(1)}].$ (4.11)
@@ 4.3 Lax connection
As in [33], we make the same ansatz for the Lax pair as in the case with constant couplings [^7] Deformation parameter and constant in front of $\sigma$-model action., namely
$\mathfrak{L}(\lambda)_{Y.B.}=K^{(0)}+\sqrt{\beta(\rho)}\left(\frac{1+\gamma^{2}}{1-\gamma^{2}}K^{(1)}+\frac{2\gamma}{1-\gamma^{2}}\star K^{(1)}\right).$ (4.12)
@@
$\mathrm{d}\star\mathrm{d}\rho-\frac{\beta^{\prime}(\rho)}{\beta(\rho)}\mathrm{d}\rho\wedge\star\mathrm{d}\rho=0,$ (4.17)
@@ 5 Outlook
A key question, also partly motivating this work, was whether it is possible to extend the techniques of integrable $D=2$ field theories to cases with dilaton potentials, i.e.
$\Box\rho=-V(\rho),$ (5.1)
with a scalar potential $V$. In this context, (pure) AdS2 dilaton-gravity with non-trivial potentials giving rise to solvable models, and even Yang-Baxter deformations thereof, have been studied [Almheiri:2014cka, Kyono:2017jtc, Kyono:2017pxs]; we stress however that these are simpler examples without coset matter coupling to $D=2$ gravity, as the ones analysed in this paper.
Neither of the deformations considered in this article did lead to such a scalar potential: the Auxiliary Field Deformation sources higher derivative modifications (3.20) and the Yang-Baxter deformations lead to a modified equation for the dilaton, which is second order in derivatives (4.17). It would therefore be compelling to understand if other kinds of integrable deformations could lead to such scalar potentials instead. This would provide new opportunities to study cosmologies in $D=2$ with the methods of integrability.
@@ C.1 Virasoro constraint from $\tau^{\infty}$ invariance of the linear system
We here follow the reasoning from [53] in order to derive the Virasoro constraints from the vanishing of the $\mathsf{K}$-coefficient in (3.24). We reformulate such condition in terms of lightcone coordinates (2.7)
@@
leads to the additional relation
@@ D.1 Review: Yang-Baxter deformation of Principal Chiral Model with spacetime dependent couplings
$\mathrm{d}\eta-\eta\frac{\mathrm{d}\rho}{\rho}=0\,.$ (D.6)
@@
$\mathfrak{L}(\lambda)=(1-c^{2}\eta^{2})\left(\frac{1}{1-\lambda^{2}}k+\frac{\lambda}{1-\lambda^{2}}\star k\right).$ (D.8)
@@
$\frac{\mathrm{d}\lambda}{\lambda}=\frac{1}{(1-c^{2}\eta^{2})(1-\lambda^{2})}\left(\left((1+\lambda^{2})(1+c^{2}\eta^{2})-4c^{2}\eta^{2}\right)\frac{\mathrm{d}\eta}{\eta}+\frac{2(\lambda^{2}-c^{2}\eta^{2})}{\lambda}\frac{\star\mathrm{d}\eta}{\eta}\right).$ (D.10)
@@
$\mathrm{d}\star\mathrm{d}\eta-\frac{\beta^{\prime}(\eta)}{\beta(\eta)}\mathrm{d}\eta\wedge\star\mathrm{d}\eta=0,$ (D.12)